# Odd-parity effect and scale-dependent viscosity in atomic quantum gases

### v1

@@ Results
For further illustration, Fig. 2(a) shows a spectral plot for the shear viscosity spectral function at finite frequency and momentum as obtained within linearized kinetic theory using a Fermi liquid two-body scattering matrix element, where we extract the Landau parameters from experimental data on the low-temperature equation of state [37]. In general, the shear viscosity spectral function is strongly peaked near zero frequency and momentum, with a frequency dependence that is well described by the hydrodynamic correlator. The trajectory $(q,\omega)=(1/R,\omega_{Q})$ in frequency-momentum space that is relevant for the damping of the quadrupole mode is indicated by the white lines for two different atom numbers.
@@
Compared to experiments in condensed matter systems, cold atomic gases offer unprecedented tunability of the universal system parameters, such as confinement, interaction strength, and particle number. Figures 2(b)-(d) show results for the anomalous enhancement of the shear viscosity evaluated at momentum $q=1/R$ and frequency $\omega=\omega_{Q}$ compared to the hydrodynamic shear viscosity (i.e., evaluated at zero momentum but finite frequency) for these different parameters. First, Fig. 2(b) shows the ratio of the shear viscosity as a function of the harmonic trapping frequency $\beta\omega_{r}$ for four interaction strengths $\ln(k_{F}a_{2})=0.39,0.49,0.56,$ and $0.63$ (corresponding to chemical potentials $\beta\mu=6.8,7.8,8.4,9.0$, respectively) at fixed atom number $N=10^{3}$ and temperature $T/T_{F}=0.1$. On the BCS side of the crossover (i.e., for $\ln(k_{F}a_{2})>0$), there is a strong enhancement of the shear viscosity over the hydrodynamic result at small frequencies. As expected and discussed above, near resonance ($\ln(k_{F}a_{2})\approx 0$) and on the BEC side ($\ln(k_{F}a_{2})<0$), the ratio quickly approaches unity, implying that there is no odd-even effect. The same vanishing of the odd-even effect on the BEC side is seen in Fig. 2(c), which shows the shear viscosity spectral function as a function interaction strength for three different values of the trap frequency $\beta\omega_{r}=0.1,0.2,$ and $0.3$, again with $N=10^{3}$ particles and temperature $T/T_{F}=0.1$. On the BCS side of the crossover and for finite confinement, the shear viscosity is enhanced due to the odd-even effect. The enhancement increases with increasing confinement and reduces to the hydrodynamic result for weak confinement. Finally, in Fig. 2(d) we examine the shear viscosity as a function of atom number at fixed interaction strength $\ln(k_{F}a_{2})=0.63$ at $T/T_{F}=0.1$ for three values of the trap frequency $\beta\omega_{r}=0.1,0.2,$ and $0.3$. Since a decreasing particle number $N$ increases the momentum as $q=1/R\sim N^{-1/4}$, we observe a strong enhancement for mesoscopic particle ensembles and a trend toward the hydrodynamic limit for large particle number.
@@ Discussion
We propose neutral quantum gases as a platform to observe and control the odd-even effect in quasiparticle lifetimes, which is an essential but to date unobserved fundamental aspect of Fermi liquid theory. Compared to electron Fermi liquids, neutral quantum gases offer three distinct advantages that will favor an experimental observation of the odd-even parity effect: First, the effect is widely tunable with interaction strength and vanishes on the BEC side of a crossover because the Fermi surface vanishes and Pauli blocking no longer provides a phase-space constraint for scattering. Second, the real-space confinement allows to directly access oscillatory transverse dynamics in real space and probe the shear viscosity from the damping of collective modes. Third, the odd-even effect depends sensitively on both trap confinement and particle number and is strongly enhanced for mesoscopic Fermi gases. Although the odd-even effect will be present for any particle number as long as the temperature is sufficiently small, our findings suggest that the onset temperature of this effect increases with decreasing particle number, providing an additional way to access this effect. Already a number of experiments have been able to create mesoscopic ensembles of fermions [50, 51, 52, 53] and would make an excellent platform for investigating this odd-even effect and its effect on the damping of the quadrupole mode.
@@ Methods
Here, $n(t,{\bf x},{\bf p})$ is the local quasiparticle distribution function of the quasiparticles with momentum ${\bf p}$ at position ${\bf x}$ and time $t$. The quasiparticle dispersion in Eq. (4) is
@@
which depends on the single-particle dispersion with atomic mass $m^{*}$ and a term that describes the mutual interaction energy from other quasiparticles. The interactions between quasiparticles in this term are parameterized by $f_{\bf p,p^{\prime}}$ and depend on $\delta n({\bf p})$, the deviation of the quasiparticle distribution function away from global equilibrium. For a circular 2D Fermi surface, we expand the interaction in angular harmonics and rescale to a dimensionless form as $F_{\bf p,p^{\prime}}=\nu f_{\bf p,p^{\prime}}=\sum_{m=-\infty}^{\infty}F_{m}e^{im(\theta_{\bf p}-\theta_{\bf p^{\prime}})}$, with $\nu=m^{*}/(\pi\hbar^{2})$ the density of states at the Fermi surface, $\theta_{\bf p}$ the polar angle of the vector ${\bf p}$ in momentum space, and $F_{m}$ the dimensionless interaction strength of the $m$-th harmonic. To determine the shear viscosity, we also include an external force ${\bf\mathcal{F}}$ in Eq. (4) induced by a gradient of a transverse velocity field,
@@
where $i,j=x,y$ are space indices, ${\bf v_{p}}=\nabla_{\bf p}\epsilon_{\bf p}$ is the quasiparticle velocity, and ${\bf V}$ is a velocity field with $\nabla_{\bf r}\cdot{\bf V}=0$. We impose a harmonic form of the velocity field with momentum ${\bf q}$ and frequency $\omega$, i.e., ${\bf V}({\bf r},t)={\bf V_{0}}e^{i{\bf q\cdot r}-i\omega t}$ with ${\bf q\cdot V_{0}}=0$.
@@
×$\times(2\pi)^{2}\delta({\bf p+q-p^{\prime}-q^{\prime}})2\pi\delta\left(\epsilon_{\bf p}+\epsilon_{\bf q}-\epsilon_{\bf p^{\prime}}-\epsilon_{\bf q^{\prime}}\right)$
×$\times\bigl{[}\left(1-n({\bf p})\right)\left(1-n({\bf q})\right)n({\bf p^{\prime}})n({\bf q^{\prime}})$
($\qquad-(\{{\bf p,q}\}\rightarrow\{{\bf p^{\prime},q^{\prime}}\})\bigr{]},$ (7)
where $\langle{\bf p^{\prime},q^{\prime}}|T|{\bf p,q}\rangle$ is the scattering $T$-matrix between states ${\bf p,q}$ and ${\bf p^{\prime},q^{\prime}}$. The linearized collision integral is then $\mathcal{L}[\psi({\bf p})]=\mathcal{I}[n_{\bf p}]/[n_{\rm FD}({\bf p})(1-n_{\rm FD}({\bf p}))]$ with the parameterization in Eq. (2). Within Fermi liquid theory, the $T$-matrix is linked to the Landau parameters $F_{m}$ [7]. In terms of the dominant Landau parameter $F_{0}$ and at low temperatures, it is a function of the scaling variable $x=\omega/(v_{F}|{\bf q-p^{\prime}}|)$,
@@
This function parametrizes repeated particle-hole scattering and is equivalent to the Lindhard function in 2D. Equation (9) holds for a well-defined Fermi surface with $\mu>0$. In the opposite limit $\mu<0$ (i.e., the BEC or high-temperature limit), other processes like particle-particle scattering will be more relevant than the ones captured by Fermi liquid theory. The supplemental material [39] discusses other forms of the scattering $T$-matrix, which give quantitatively similar results. The dominant Landau parameter $F_{0}$ is connected to the inverse compressibility as $\kappa^{-1}=(1+F_{0})\kappa_{0}^{-1}$ with $\kappa_{0}^{-1}$ the inverse compressibility of a non-interacting Fermi gas. The compressibility has been previously measured experimentally over the entire BCS-BEC crossover [54, 37, 36], which gives the Landau parameter $F_{0}$ in terms of the interaction parameter $\ln(k_{F}a_{2})$. The fit is discussed in the supplemental material [39].
@@
Acknowledgments This work is supported by Vetenskapsrådet (Grant No. 2020-04239) (JH). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Linköping University partially funded by the Swedish Research Council through grant agreement no. 2022-06725. JM is partially supported by the Provincia Autonoma di Trento.
Author contributions All authors contributed to conception, computations, and preparation of the manuscript.
@@
Correspondence and requests for materials should be addressed to Jeff Maki.

### v2

@@ Results
For further illustration, Fig. 2(a) shows a spectral plot of the shear viscosity spectral function at finite frequency and momentum as obtained using a Fermi liquid two-body scattering matrix element, where we extract the Landau parameters from experimental data on the low-temperature equation of state [37]. In general, the shear viscosity spectral function is related to the retarded correlation function of the off-diagonal elements of the stress tensor [48], the poles of which give the transverse collective modes. However, in our case the transverse sound pole is diffusive, as the transverse sound mode only becomes real when $F_{1}>1$ [49]. At zero momentum and finite frequency, we find that the spectral function has a Lorentzian shape that describes the standard hydrodynamic-to-collisionless crossover [42, 43, 44]. At finite momentum, the shear viscosity spectral function vanishes at small frequencies, which reflects the fact that the fluid responds instantaneously to the transverse velocity field. At large frequencies the shear viscosity is known to decay as $\omega^{-2}$ in the collisionless limit [32]. Hence one observes a local maximum in the shear viscosity spectral function at finite momentum. The trajectory $(q,\omega)=(1/R,\omega_{Q})$ that is traced out in frequency-momentum space relevant for the damping of the quadrupole mode is indicated by the white lines in Fig. 2(a) for two different atom numbers.
@@
Compared to experiments in condensed matter systems, cold atomic gases offer unprecedented tunability of the universal system parameters, such as confinement, interaction strength, and particle number. Figures 2(b)-(d) show results for the anomalous enhancement of the shear viscosity for these different parameters evaluated at momentum $q=1/R$ and frequency $\omega=\omega_{Q}$ compared to the hydrodynamic shear viscosity (i.e., evaluated at zero momentum but finite frequency). First, Fig. 2(b) shows the ratio of the shear viscosity as a function of the harmonic trapping frequency $\beta\omega_{r}$ for four interaction strengths $\ln(k_{F}a_{2})=0.39,0.49,0.56,$ and $0.63$ (corresponding to chemical potentials $\beta\mu=6.8,7.8,8.4,9.0$, respectively) at fixed atom number $N=10^{3}$ and temperature $T/T_{F}=0.1$. On the BCS side of the crossover (i.e., for $\ln(k_{F}a_{2})>0$), there is a strong enhancement of the shear viscosity over the hydrodynamic result at small frequencies. As expected we observe hydrodynamic transport in the strongly interacting limit ($\ln(k_{F}a_{2})\approx 0$) implying there is no odd-even effect. This trend continues on the BEC side ($\ln(k_{F}a_{2})<0$) where the odd- and even-parity relaxation rates are of the same order of magnitude. This can be seen in Fig. 2(c), which shows the shear viscosity spectral function as a function interaction strength for three different values of the trap frequency $\beta\omega_{r}=0.1,0.2,$ and $0.3$, again with $N=10^{3}$ particles and temperature $T/T_{F}=0.1$. On the BCS side of the crossover and for finite confinement, the shear viscosity is enhanced due to the odd-even effect. The enhancement increases with increasing confinement and reduces to the hydrodynamic result for weak confinement. Finally, in Fig. 2(d) we examine the shear viscosity as a function of atom number at fixed interaction strength $\ln(k_{F}a_{2})=0.63$ at $T/T_{F}=0.1$ for three values of the trap frequency $\beta\omega_{r}=0.1,0.2,$ and $0.3$. Since a decreasing particle number $N$ increases the momentum as $q=1/R\sim N^{-1/4}$, we observe a strong enhancement for mesoscopic particle ensembles and a trend toward the hydrodynamic limit for large particle number.
Comparing these results with the criterion in Eq. (4), we indeed find that it accurately predicts the onset of the tomographic regime. As an example, for parameters where we observe an anomalous shear viscosity—such as for $\beta\omega_{r}=0.1$, $\ln(k_{F}a)\approx 10$ and $N=10^{3}$—we find $q\xi\approx 3.2$. By contrast, near resonance with $\ln(k_{F}a)\approx 1$ with the same $\beta\omega_{r}$ and $N$, we find $q\xi\approx 0.05$.
@@ Discussion
We propose neutral quantum gases as a platform to observe and control the odd-even effect in quasiparticle lifetimes, which is an essential but to date unobserved fundamental aspect of Fermi liquid theory. Compared to electron Fermi liquids, neutral quantum gases offer three distinct advantages that will favor an experimental observation of the odd-even parity effect: First, the effect is widely tunable with interaction strength and vanishes on the BEC side of the crossover because the Fermi surface vanishes and Pauli blocking no longer provides a phase-space constraint for scattering. Second, the real-space confinement allows to directly access oscillatory transverse dynamics in real space and probe the shear viscosity through the damping of collective modes. While we here discuss the quadrupole mode damping in detail, in principle, one could also excite other collective modes, like the breathing mode, which is isotropic and will not depend sensitively on the odd-even effect, as a benchmark to examine the anomalous behaviour of the shear viscosity. Third, the odd-even effect depends sensitively on both trap confinement and particle number and is strongly enhanced for mesoscopic Fermi gases. Although the odd-even effect will be present for any particle number as long as the temperature is sufficiently small, our findings suggest that the onset temperature of this effect increases with decreasing particle number, providing an additional way to access this effect. Already a number of experiments have been able to create mesoscopic ensembles of fermions [50, 51, 52, 53] and would make an excellent platform for investigating this odd-even effect and its effect on the damping of the quadrupole mode.
@@ Methods
Here, $n({\bf p})=n(t,{\bf x},{\bf p})$ is the local distribution function of quasiparticles at time $t$ and position $\bf r$} with momentum $\bf p$} (we omit the space-time dependence in the following). The quasiparticle dispersion in Eq. (5) is
@@
[…]
@@
where $i,j=x,y$ are space indices, ${\bf v_{p}}=\nabla_{\bf p}\epsilon_{\bf p}$ is the quasiparticle velocity, and $\bf V$} is a velocity field with $\nabla_{\bf r}\cdot{\bf V}=0$. We impose a harmonic form of the velocity field with momentum $\bf q$} and frequency $\omega$, i.e., ${\bf V}={\bf V_{0}}e^{-i\omega t+i{\bf q\cdot r}}$ with $\bf V_{0}$} a constant vector satisfying ${\bf q\cdot V_{0}}=0$.
@@
$\times(2\pi)^{2}\delta({\bf p+q-p^{\prime}-q^{\prime}})2\pi\delta\left(\epsilon_{\bf p}+\epsilon_{\bf q}-\epsilon_{\bf p^{\prime}}-\epsilon_{\bf q^{\prime}}\right)$
$\times\bigl{[}\left(1-n({\bf p})\right)\left(1-n({\bf q})\right)n({\bf p^{\prime}})n({\bf q^{\prime}})$
$\qquad-(\{{\bf p,q}\}\rightarrow\{{\bf p^{\prime},q^{\prime}}\})\bigr{]},$ (8)
where $\langle{\bf p^{\prime},q^{\prime}}|T|{\bf p,q}\rangle$ is the scattering $T$-matrix between states $\bf p,q$} and $\bf p^{\prime},q^{\prime}$}. The linearized collision integral is then $\mathcal{L}[\psi({\bf p})]=\mathcal{I}[n_{\bf p}]/[n_{\rm FD}({\bf p})(1-n_{\rm FD}({\bf p}))]$ with the parameterization of Eq. (2). Within Fermi liquid theory, the $T$-matrix is linked to the Landau parameters $F_{m}$ [7]: In terms of the dominant Landau parameter $F_{0}$ and at low temperatures, it is a function of the scaling variable $x=\omega/(v_{F}|{\bf q-p^{\prime}}|)$,
@@
This function parametrizes repeated particle-hole scattering and is equivalent to the Lindhard function in 2D. Equation (10) holds for a well-defined Fermi surface with $\mu>0$. In the opposite limit $\mu<0$ (i.e., the BEC or the high-temperature limit), other processes like particle-particle scattering will be more relevant than the ones captured by Fermi liquid theory. For good measure, Supplementary Note I compares and contrasts other forms of the scattering $T$-matrix, which give quantitatively similar results. The dominant Landau parameter $F_{0}$ is connected to the inverse compressibility as $\kappa^{-1}=(1+F_{0})\kappa_{0}^{-1}$ with $\kappa_{0}^{-1}$ the inverse compressibility of a non-interacting Fermi gas. This compressibility has been previously measured experimentally over the entire BCS-BEC crossover [54, 37, 36], which gives the Landau parameter $F_{0}$ in terms of the interaction parameter $\ln(k_{F}a_{2})$. The fit is discussed in the Supplementary Note III.
@@
Acknowledgments JM is partially supported by the Provincia Autonoma di Trento. This work is supported by Vetenskapsrådet (Grant Nos. 2020-04239 and 2024-04485), the Olle Engkvist Foundation (Grant No. 233-0339), the Knut and Alice Wallenberg Foundation (Grant No. KAW 2024.0129), and Nordita. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Linköping University partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Author contributions JM, UG, and JH all contributed to equally to the conception, computations, and preparation of the manuscript.
@@
Correspondence should be addressed to any of the authors.
Code availability requests should be addressed to any of the authors.
Data availability The raw data for the figures are available upon request addressed to any of the authors.